OpenStudy (anonymous):

Solve for v: v^2+7v+10=0

OpenStudy (anonymous):

v^2+5v+2v+10=0 Now carry on

OpenStudy (anonymous):

So you would add 5v and 2v and have the answer?

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

v(v+5)+2(v+5)=0 (v+2)(v5)=0 so v=-2,-5

OpenStudy (anonymous):

v+5 sorry!

OpenStudy (anonymous):

So the answer would be v^2+7v+10=0?

OpenStudy (anonymous):

see my calculation